OpenStudy (anonymous):

find the polynomial that has the root at -3 then 4 with a multiplicity of 2?

OpenStudy (anonymous):

anyone ?

zepdrix (zepdrix):

Do you understand how to write roots when constructing a polynomial? A root of -3 can be written this way,\[\Large x=-3 \qquad\qquad\to\qquad\qquad 0=(x+3)\] We also have a root at 4,\[\Large x=4 \qquad\qquad\to\qquad\qquad 0=(x-4)\]We'll multiply our roots together to form a polynomial,\[\Large 0=(x+3)(x-4)\]They told us that the second root listed has a multiplicity of 2. I think that means the root occurs twice, so we would multiply by that root again,\[\Large 0=(x+3)(x-4)(x-4)\]

zepdrix (zepdrix):

To find our polynomial, we would multiply this all out.

zepdrix (zepdrix):

You leftttttt +_+ so rude! lol

OpenStudy (anonymous):

im so sorry lol but thank you so much

zepdrix (zepdrix):

:3

OpenStudy (anonymous):

is that the answer ?

zepdrix (zepdrix):

Multiplying the factors together that give us our roots:\[\Large 0=(x+3)(x-4)(x-4)\]No, for a final answer, you would want to expand out the brackets :o Just a bit of multiplication from here.

zepdrix (zepdrix):

Understand how to multiply those? :o

OpenStudy (anonymous):

kindaaaaaa nott reaaaaallly LOL :p

zepdrix (zepdrix):

you -_-

zepdrix (zepdrix):

Wait I'm so confused... you asked this question like a WEEK ago XD How is it still relevant? :X

OpenStudy (anonymous):

hahah :P whhhaaatttt im sowwwy u can aske me to write a 5 million word essay and i could do that Math im likd WHATTTTT :p

zepdrix (zepdrix):

Ugh writing T.T I struggle just reading a large paragraph.. English was so brutal... :c

OpenStudy (anonymous):

because im going to retake a test that i failed in 3 min

zepdrix (zepdrix):

oh my :o

OpenStudy (anonymous):

see I LOVE ENGLISH everyone has there talents gosh :)

zepdrix (zepdrix):

true story c:

OpenStudy (anonymous):

i know :0

OpenStudy (anonymous):

its only 5 questions though :)

zepdrix (zepdrix):